OpenStudy (anonymous):

How do I write y=2x-8 in standard form?

OpenStudy (anonymous):

Standard form of a line is Ax+By=C

OpenStudy (anonymous):

\[y = 2 (x-4), -2 x+y+8 = 0\]

OpenStudy (anonymous):

So you would take y=2x-8 and subtract 2x on both sides. -2x - y = -8